OpenStudy (anonymous):

square root 2 add square root 18 add square root 32

OpenStudy (anonymous):

1.41+4.24+5.65=11.3

jhonyy9 (jhonyy9):

sqrt2 +sqrt18 +sqrt32 = sqrt2 +3sqrt2 +4sqrt2 = 8sqrt2 hope this will help you